One number is 5 less than three-fourths the other number. Their Sum is 79. Find the two numbers
Answer by fractalier(6550)

Call the numbers x and y.
x = (3/4)y - 5 and
x + y = 79 which leads to x = 79 - y
Since they both equal x we can set them equal and get
(3/4)y - 5 = 79 - y
Add y to both sides and add five to both sides and get
1.75y = 84
Now divide by 1.75 and get
y = 48
x = 31